OpenStudy (anonymous):

If f(x) = 7x - 1, what does f(12) represent?

OpenStudy (anonymous):

The value if x when (7x - 1) = 12 The value of (7x - 1) when x = 12 The value of 12 f(x) when x = 12 The value of f(x) when (7x - 1) = 12

OpenStudy (anonymous):

the value of 7x-1 when x=12

OpenStudy (anonymous):

Are you sure?

OpenStudy (anonymous):

yeah, you plug the 12 in the x in the 7x-1

OpenStudy (anonymous):

DON'T WORRY i'm studying on functions too hey do you know how to show that 2 functions are inverses of each other?

OpenStudy (anonymous):

Oh okay :) Thank you! And not really :( I'm super sorry

OpenStudy (anonymous):

it's cooooll...